OpenStudy (anonymous):

Which pair of points represents the solution set of this system of equations? y = x2 − 6x + 8 2y + x = 4 a (2, 1) and (4, 0) b (1.5, 1.25) and (4, 0) c (0, 0) and (5, -5) d (-1.25, -2.75) and (1.25, 3) e There is no real solution. NextReset

OpenStudy (anonymous):

I think its a or c

OpenStudy (anonymous):

Can be solved with substitution: \[2y+x=4 \rightarrow y=(4-x)/2.\] \[\rightarrow(4-x)/2=x^2-6x+8\] Try simplifying that. If you have trouble factoring then just use the quadratic formula.